# SECURA: Sigmoid-Enhanced CUR Decomposition with Uninterrupted Retention and Low-Rank Adaptation in Large Language Models

### v3

@@ Abstract
With the rapid development of large language models (LLMs), fully fine-tuning(FT) these models has become less feasible due to high computational demands. Moreover, FT also increases the risk of catastrophic forgetting. As an alternative, Low-Rank Adaptation (LoRA) has been proposed. By fine-tuning only a small subset of parameters, LoRA achieves performance similar to FT while significantly reducing resource requirements. However, since LoRA inherits FT’s design, the issue of catastrophic forgetting still remains. To address these limitations, we propose SECURA: Sigmoid-Enhanced CUR Decomposition LoRA, a novel PEFT variant designed to mitigate catastrophic forgetting while improving fine-tuning performance. Our method introduces a new normalization technique, SigNorm, to enhance parameter retention and improve overall performance. SECURA has been evaluated on a diverse range of tasks, including mathematical problem-solving (GSM8K), challenging question-answering (CNNDM), translation (NewsDE), and complex multiple-choice reasoning (LogiQA). Experimental results demonstrate it achieves an average fine-tuning improvement of 3.59% across 4 MCQ sphere tasks and a 2.51% improvement in 5 QA sphere tasks across Gemma2 2b, Qwen2 1.5b, Qwen 2 7b,Llama3 8b,Llama3.1 8b compared to DoRA. Additionally, SECURA demonstrates superior knowledge retention capabilities, achieves state-of-the-art in 16 continual learning tests keep more than 70% accuracy on LLMs basic knowledge compared with Experience Replay(ER), sequential learning(SEQ),EWC, I-LoRA and CUR-LoRA.
@@ 1 Introduction
However, LoRA remains an approximation of FT, meaning it hardly match or exceed the performance of full fine-tuning. Furthermore, it also inherits FT’s critical issue: catastrophic forgetting (McCloskey and Cohen, 1989) . While freezing the base model weights alleviates this problem to some extent, the addition of LoRA matrices reintroduces the issue. Additionally, fine-tuning fewer parameters, while reducing computational cost, increases the risk of catastrophic forgetting before the model finds an optimal solution. Although Weight normalization techniques (Salimans and Kingma, 2016) can be used to mitigate these issues but have not provided a comprehensive solution. Existing LoRA variants primarily focus on improving specific aspects: enhancing performance (e.g. DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2023)), accelerating convergence (e.g., PiSSA (Meng et al., 2024)), or only reducing catastrophic forgetting (e.g. I-LoRA (Ren et al., 2024), CUR-LoRA (Fawi, 2024)). However, few approaches simultaneously address uninterrupted retention of knowledge and keep the performance.
@@
  • SigNorm Normalization: To balance model stability and performance, we designed SigNorm—a normalization method using Sigmoid’s gradual transition property. It dynamically adjusts parameters during training to prevent catastrophic forgetting by prioritizing the collective impact of central parameter changes over extreme values, eliminating the need for historical data or prior outputs. This ensures balanced weight updates while maintaining adaptability.
By integrating SigNorm and CABR decomposition, SECURA effectively addresses memory retention challenges in continual learning while surpassing existing LoRA methods and their variants in accuracy under comparable computational budgets. Experiments on 5 LLMs (Gemma2 2B, Qwen2 1.5B/7B, Llama3 8B, Llama3.1 8B) demonstrate SECURA’s superiority: it achieves 3.63% average improvement on 4 multiple-choice (MCQ) tasks and 2.56% on 5 question-answering (QA) tasks over baseline LoRA, even rivaling advanced variants like DoRA. For knowledge retention, SECURA maintains over 70% accuracy on fine-tuned test sets, outperforming state-of-the-art methods such as Experience Replay (ER) (Fedus et al., 2020), sequential learning (SEQ) (Sutskever et al., 2014), EWC (Aich, 2021),I-LoRA (Ren et al., 2024) and CUR-LoRA (Fawi, 2024). These results highlight SECURA’s dual strengths in task adaptation and preserving pre-trained knowledge.
@@ 2 Background
Figure 1: SigNorm Update method: The figure illustrates the process flow of the SigNorm Normalization algorithm. Shows the steps of starting with the fusion of the former weight matrix with the CABR module and moving through the ratio loss matrix calculation, the normalization and sigmoid steps, followed by the final limiting of the ratio loss matrix.
@@ 3 Methodology
Our proposed SECURA method is based on three key components: CABR-LoRA inverse decomposition fine-tuning, the SigNorm Normalization algorithm, and two types of SECURA Merge method. Drawing inspiration from dynamic network pruning and CUR-decomposition, we hypothesize that parameters with higher norms in the original model are more important, as they store more critical knowledge. By constraining these important parameters and focusing on fine-tuning the less important ones, SECURA can achieve remarkable performance while significantly mitigating catastrophic forgetting. Appendix B shows all intuitive equation derivation.
Figure 2: Gradient Analysis: The gradient change during training shows LoRA has higher gradient fluctuations, indicating greater parameter reshaping, which may increase the risk of catastrophic forgetting. In contrast, SECURA shows lower gradient changes, avoided excessive drift, suggesting more stable parameter updates, improving catastrophic forgetting mitigation and ability of finding a better optimum for the model’s parameters.
@@ 3.1 CABR-LoRA Inverse Decomposition Initialization
CABR-LoRA is based on CUR-LoRA, utilizing CUR decomposition to break down the weight matrices. As shown in Figure 1, CABR-LoRA is initialized using a inverse low-rank decomposition of the $U$ matrix. This allows the "less important" columns and rows to remain smaller while maintaining a higher-dimensional space for fine-tuning SECURA. The formula for initialize of SECURA is:
@@
Where $\mathcal{W}_{\text{CABRWeight}}$ is the initialized CABR-LoRA weight, $C$ and $R$ represent the less important columns and rows, while $A_{\text{initial}}$ and $B_{\text{zeros}}$ are the trainable matrices with weights $\mathcal{W}_{\text{A}}$ and $\mathcal{W}_{\text{B}}$.
CUR-LoRA reduces the number of trainable parameters to locate lest important $\boldsymbol{C}$ and $\boldsymbol{R}$, but this often leads to the loss of important parameters, causing performance degradation. To address this, we introduce CABR decomposition, which decomposes the $U$ matrix, into two matrices: $\mathcal{W}_{\text{A}}$ and $\mathcal{W}_{\text{B}}$, which shapes are $r\times m$ and $m\times r$. Respectively here $m>r$. $\mathcal{W}_{\text{A}}$ is initialized via SVD decomposition, while $\mathcal{W}_{\text{B}}$ is initialized with zeros. By initializing $\mathcal{W}_{A}$ with SVD decomposition, CABR-LoRA effectively preserves prior knowledge. The initialization of $\mathcal{W}_{B}$ with zeros ensures that this matrix will fine-tune from the base LLM without altering the critical knowledge. The decomposition follows the formulas below:
$\mathcal{W}_{\text{frozen}}=USV^{T}.$ (2)
$\mathcal{W}_{A}=\mathcal{U}_{\text{[:r,:]}}\cdot\mathcal{S}_{\text{[:,:]}}\cdot V_{\text{[:m,:]}}^{T}$ (3)
$\mathcal{W}_{B}=\mathcal{M}_{\text{Zeros}},\quad\text{where}\quad\mathcal{M}_{\text{Zeros}}\in\mathbb{R}^{m\times r}.$ (4)
3.2 SigNorm Normalization Algorithm
While CABR decomposition provides a higher dimension of fine-tuning, only using CABR-LoRA alone does not fully address catastrophic forgetting, which is crucial for performance in lifelong learning. To tackle this, we introduce the SigNorm Normalization algorithm, inspired by dynamic network pruning. This algorithm adjusts the weight updates by controlling the change magnitude between the former and new weights, provide a stable parameter update (figure 2).
The core idea is to employ a Sigmoid-based dynamic scaling mechanism to regulate weight updates. This method constructs a weight pruning matrix whose values are constrained within the range $[1,2]$. The Sigmoid function primarily emphasizes central value adjustments, effectively suppressing extreme variations. We can allow SECURA deal with whole matrix’s changing tendency while limiting the extreme parameters’ effect. At the same time, restricting the update amplitude of lower-norm columns and rows ensures that important parameters(with higher norms)’ magnitude experience minimal change, while less important parameters (with smaller norms) undergo more significant updates. The former weight matrix is denoted as $W_{\text{former}}$, and the SECURA’s matrices as $A$ and $B$.
@@ 3.2 SigNorm Normalization Algorithm
$\mathcal{W}_{\text{SECURAMerged}}=CABR+\mathcal{W}_{\text{former}}.$ (5)
Next, we calculate the Magnitude Loss Matrix $\mathcal{M}_{\text{Mag}}$, which captures the relative magnitude change between the new and former weights. Here $\epsilon$ is a small value to avoid division by too small parameters:
$\mathcal{M}_{\text{Mag}}=\left|\frac{\mathcal{W}_{\text{SECURAMerged}}}{\mathcal{W}_{\text{former}}+\epsilon}\right|.$ (6)
@@
Finally, we apply the Sigmoid function to the normalized matrix to produce the SigNorm values and get the final restring matrix $\mathcal{M}_{\text{Res}}$:
@@
Once we have weight pruning matrix $\mathcal{M}_{\text{Res}}$ , the megred matrix $W_{\text{SECURAMerged}}$, will then adjust the update process by dividing the SigNorm pruning-weight matrix $\mathcal{M}_{\text{Res}}$:
@@
3.3 Merge with Basic Model:
Limiting parameter growth alone is not the optimal solution. In our early experiments, we found that using only the methods from Sections 3.1 and 3.2 led to worse performance—not only in basic performance but also in mitigating catastrophic forgetting. This issue was identified through insights from delayed updates (Kosta et al., 2020) and Target Network (Gao et al., 2017). We realized that while basic parameters should not remain static during long-term training, they must be updated to prevent performance from stagnating or even deteriorating.
If the basic parameters are kept unchanged, two cycle emerges: initially, unimportant parameters are updated, while important ones are constrained. However, in the second cycle, important parameters become smaller than unimportant ones, and the cycle worsens as important parameters begin to change more rapidly, causing further catastrophic forgetting and worse performance.
To address this issue, we propose two merging methods:
Merge with basic weight(M1 method):
This approach requires only one SECURA training cost but does not keep the basic LLM weights unchanged. The update rule is as follows:
$\mathcal{W}_{\text{Newformer}}=CABR+\mathcal{W}_{\text{former}}.$ (10)
Merge with SECURA(M2 method):
This method uses two identical SECURA models but freezes the base LLM weights. An additional calculation is performed to obtain the final updated weights The update rule is as follows:
$\mathcal{W}_{\text{former}}=CA_{\text{former}}B_{\text{former}}R+\mathcal{W}_{\text{basic}}.$ (11)
The merge process is described by the following formulas:
$A_{\text{Newformer}}=A_{\text{train}}.$ (12)
$B_{\text{Newformer}}=B_{\text{former}}+B_{\text{train}}.$ (13)
In both M1 and M2, the new parameters of $B_{\text{former}}$ must be reset to zero after each update, while the matrix remains unchanged. This helps the LLM learn new knowledge effectively.
For the two different merging methods, M1 achieves better fine-tuning performance but at the cost of some knowledge retention, while M2 preserves more knowledge but sacrifices fine-tuning effectiveness due to its shallower merging. Based on our ablation study (4), we recommend M2 for tasks emphasizing knowledge retention (e.g., medical and economics tasks) and M1 for tasks prioritizing fine-tuning performance (e.g., mathematics and programming tasks). Overall, M1 strikes a better balance, maintaining strong knowledge retention while achieving superior fine-tuning results.
Figure 3: Comparison of Weight SVD Norm across different methods: The "Basic" method represents the original LLM weights without fine-tuning. After 2000 training steps, DoRA shows a 0.517 decrease in norm, indicating some loss of information. SECURA preserves more knowledge with a minor 0.061 change, while LoRA shows a significant 3.076 increase, reflecting greater weight modification.
@@ 4.1 Experimental Setup
To evaluate catastrophic forgetting, we combined three QA datasets-ARC-Challenge(ARC-C) (Clark et al., 2018), GSM8K (Cobbe et al., 2021), Multipl-E (Cassano et al., 2022)-as the test set , which were used training/fine-tuning LLMs we chosen. This fusion dataset will be referred to as F-task in following parts. Further details can be found in the appendix C.
@@
Experiment 3 (Catastrophic Forgetting with Knowledge retention method): Benchmarks against continual learning methods at 1e-5 learning rate with 32K steps 16 task sequential training(each training 2k steps) every PEFT based training method will fusion itself to frozen basic weights(e.g. CUR-LoRA, I-LoRA, SECURA) after one training task is done, measuring remain knowledge by using the F-task set and measured by BLEU, ROUGE-1, ROUGE-L.
Note: All main experiments use the M1 merge method. For hyperparameter details, refer to the appendix D.

### v4

@@ Abstract
With the rapid development of large language models (LLMs), fully fine-tuning(FT) these models is becoming increasingly infeasible due to high computational demands. Moreover, FT also increases the risk of catastrophic forgetting. As an alternative, Low-Rank Adaptation (LoRA) has been proposed. By fine-tuning only a small subset of parameters, LoRA achieves performance similar to FT while significantly reducing resource requirements. However, since LoRA inherits FT’s design, the issue of catastrophic forgetting still remains. To address these limitations, we propose SECURA: Sigmoid-Enhanced CUR Decomposition LoRA, a novel PEFT variant designed to mitigate catastrophic forgetting while improving fine-tuning performance. Our method introduces a novel normalization technique, Sigmoid based Magnitude Norm(S-MagNorm), which enhances parameter retention and improve fine-tuning performance. SECURA has been evaluated on a diverse range of tasks, including mathematical problem-solving (GSM8K), complex question-answering (CNNDM), translation (NewsDE), and complex multiple-choice reasoning (LogiQA). Experimental results demonstrate that it achieves an average fine-tuning improvement of 3.59% across 4 MCQ sphere tasks and a 2.51% improvement in 5 QA sphere tasks across Gemma2 2b, Qwen2 1.5b, Qwen2 7b,Llama3 8b,Llama3.1 8b compared to DoRA. Additionally, SECURA demonstrates superior knowledge retention capabilities, achieves state-of-the-art performance in 16 continual learning tests maintaining more than 70% accuracy on LLMs’ basic knowledge compared with Experience Replay(ER), sequential learning(SEQ),EWC, I-LoRA and CUR-LoRA.
@@ 1 Introduction
However, LoRA remains an approximation of FT, meaning it can hardly match or exceed the performance of full fine-tuning. Furthermore, it also inherits FT’s critical issue: catastrophic forgetting (McCloskey and Cohen, 1989) . While freezing the base model weights alleviates this problem to some extent, the addition of LoRA matrices reintroduces the issue. Additionally, fine-tuning fewer parameters, while reducing computational cost, increases the risk of catastrophic forgetting before the model finds an optimal solution. Although weight normalization techniques (Salimans and Kingma, 2016) can be used to mitigate these issues, they do not provide a comprehensive solution. Existing LoRA variants primarily focus on improving specific aspects: enhancing performance (e.g. DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2023)), accelerating convergence (e.g., PiSSA (Meng et al., 2024)), or only reducing catastrophic forgetting (e.g. I-LoRA (Ren et al., 2024), CUR-LoRA (Fawi, 2024)). However, few approaches simultaneously address uninterrupted retention of knowledge and keep the performance.
@@
  • S-MagNorm Normalization: To balance model stability and performance, we designed S-MagNorm—a normalization method using Sigmoid’s gradual transition property. It dynamically adjusts parameters during training to prevent catastrophic forgetting by prioritizing the collective impact of central parameter changes over extreme values, eliminating the need for historical data or prior outputs. This ensures balanced weight updates while maintaining adaptability.
By integrating S-MagNorm and CABR decomposition, SECURA effectively addresses memory retention challenges in continual learning while surpassing existing LoRA methods and their variants in accuracy under comparable computational budgets. Experiments on 5 LLMs (Gemma2 2B, Qwen2 1.5B/7B, Llama3 8B, Llama3.1 8B) demonstrate SECURA’s superiority: it achieves 3.63% average improvement on 4 multiple-choice (MCQ) tasks and 2.56% on 5 question-answering (QA) tasks over baseline LoRA, even rivaling advanced variants like DoRA. For knowledge retention, SECURA maintains over 70% accuracy on fine-tuned test sets, outperforming state-of-the-art methods such as Experience Replay (ER) (Fedus et al., 2020), sequential learning (SEQ) (Sutskever et al., 2014), EWC (Aich, 2021),I-LoRA (Ren et al., 2024) and CUR-LoRA (Fawi, 2024). These results highlight SECURA’s dual strengths in task adaptation and preserving pre-trained knowledge.
@@ 2 Background
Figure 1: S-MagNorm Update method: The figure illustrates the process flow of the S-MagNorm Normalization algorithm. Shows the steps of starting with the fusion of the former weight matrix with the CABR module and moving through the ratio loss matrix calculation, the normalization and sigmoid steps, followed by the final limiting of the ratio loss matrix.
@@ 3 Methodology
Our proposed SECURA method is based on three key components: CABR-LoRA inverse decomposition fine-tuning, the S-MagNorm Normalization algorithm, and two types of SECURA Merge method. Drawing inspiration from dynamic network pruning and CUR-decomposition, we hypothesize that parameters with higher norms in the original model are more important, as they store more critical knowledge. By constraining these important parameters and focusing on fine-tuning the less important ones, SECURA can achieve remarkable performance while significantly mitigating catastrophic forgetting. Appendix B shows all intuitive equation derivation.
Figure 2: Gradient Analysis: The gradient variations during training shows that LoRA exhibits higher fluctuations, indicating greater parameter reshaping, which may increase the risk of catastrophic forgetting. In contrast, SECURA (CABR-LoRA + S-MagNorm) and CABR-LoRA Only shows lower gradient changes, avoided excessive drift, suggesting more stable parameter updates, improving catastrophic forgetting mitigation and ability of finding a better optimum for the model’s parameters. Additionally, the range and variance comparsion shows the essential role of S-MagNorm. Hyperparameter settings are detailed in Appendix D.
@@ 3.1 CABR-LoRA Inverse Decomposition Initialization
CABR-LoRA is based on CUR-LoRA, utilizing CUR decomposition to break down the weight matrices. As shown in Figure 1, CABR-LoRA is initialized using a inverse low-rank decomposition of the $U$ matrix. This allows the "less important" columns and rows to remain smaller while maintaining a higher-dimensional space for fine-tuning SECURA. Where $\mathcal{W}_{\text{CABRWeight}}$ is the initialized CABR-LoRA weight, $C$ and $R$ represent the less important columns and rows, while $A_{\text{initial}}$ and $B_{\text{zeros}}$ are the trainable matrices with weights $\mathcal{W}_{\text{A}}$ and $\mathcal{W}_{\text{B}}$. The initialization formula for SECURA is given as follows:
@@
CUR-LoRA reduces the number of trainable parameters to locate least important $\boldsymbol{C}$ and $\boldsymbol{R}$, but this often leads to the loss of important parameters, causing performance degradation. To address this, we introduce CABR decomposition, which decomposes the $U$ matrix, into two matrices: $\mathcal{W}_{\text{A}}$ and $\mathcal{W}_{\text{B}}$, which shapes are $r\times m$ and $m\times r$. Respectively here $m>r$. $\mathcal{W}_{\text{A}}$ is initialized via SVD decomposition, while $\mathcal{W}_{\text{B}}$ is initialized with zeros. By initializing $\mathcal{W}_{A}$ with SVD decomposition, CABR-LoRA effectively preserves prior knowledge. The initialization of $\mathcal{W}_{B}$ with zeros ensures that this matrix will fine-tune from the base LLM without altering the critical knowledge. The decomposition follows the formulas below:
$\mathcal{W}_{\text{frozen}}=\mathcal{U}_{\text{SVD}}\mathcal{S}_{\text{SVD}}\mathcal{V}_{\text{SVD}}^{\top}.$ (2)
$\mathcal{W}_{A}=\mathcal{U}_{\text{SVD}}^{(r,:)}\mathcal{S}_{\text{SVD}}\mathcal{V}_{\text{SVD}}^{(:,m)\top}.$ (3)
$\mathcal{W}_{B}=\mathbf{0}_{m\times r},\quad\text{where}\quad\mathbf{0}_{m\times r}\text{ is a zero matrix}.$ (4)
3.2 S-MagNorm Normalization Algorithm
While CABR decomposition provides a higher dimension of fine-tuning, only using CABR-LoRA alone does not fully address catastrophic forgetting, which is crucial for performance in lifelong learning. To tackle this, we introduce the S-MagNorm Normalization algorithm, inspired by dynamic network pruning. This algorithm adjusts the weight updates by controlling the change magnitude between the former and new weights, provide a stable parameter update (figure 2).
The core idea is to employ a Sigmoid-based dynamic scaling mechanism to regulate weight updates. This method constructs a weight pruning matrix whose values are constrained within the range $[1,2]$ to achieve the task of weighted mean between basic matrix weight and updated matrix weight. The Sigmoid function primarily emphasizes central value adjustments, effectively suppressing extreme variations. We can allow SECURA deal with whole matrix’s changing tendency while limiting the extreme parameters’ effect. Simultaneously, restricting the update amplitude of lower-norm columns and rows ensures that important parameters(with higher norms)’ magnitude experience minimal change, while less important parameters (with smaller norms) undergo more significant updates. The base weight matrix is denoted as $W_{\text{base}}$, and the SECURA’s matrices as $A$ and $B$.
@@ 3.2 S-MagNorm Normalization Algorithm
$\mathcal{W}_{\text{SECURAMerged}}=CABR+\mathcal{W}_{\text{base}}.$ (5)
Next, we calculate the Magnitude Loss Matrix $\mathcal{M}_{\text{Mag}}$, which captures the relative magnitude change between the new and base weights. Here $\epsilon$ is a small value to avoid division by too small parameters:
$\mathcal{M}_{\text{Mag}}=\left|\frac{\mathcal{W}_{\text{SECURAMerged}}}{\mathcal{W}_{\text{base}}+\epsilon}\right|.$ (6)
@@
Finally, we apply the Sigmoid function to the normalized matrix to produce the S-MagNorm values and get the final restring matrix $\mathcal{M}_{\text{Res}}$:
@@
Once we have weight pruning matrix $\mathcal{M}_{\text{Res}}$ , the mergred matrix $W_{\text{SECURAMerged}}$, will then adjust the update process by dividing the S-MagNorm pruning-weight matrix $\mathcal{M}_{\text{Res}}$:
@@
3.3 Merge with Basic Model
Limiting parameter growth alone is not optimal. Our early experiments showed that using only CABR-LoRA and S-MagNorm degraded both base performance and catastrophic forgetting mitigation. Drawing insights from delayed updates (Kosta et al., 2020) and Target Network (Gao et al., 2017), we recognized that base parameters require controlled updates to prevent performance stagnation.
When freezing base parameters, a destructive two-phase cycle emerges: 1) Unimportant parameters update freely while important ones remain constrained. 2) Important parameters become smaller than unimportant ones in subsequent cycles, accelerating catastrophic forgetting as critical knowledge degrades.
We resolve this through two merging strategies:
M1: Direct Weight Merge
Updates base LLM weights with one SECURA training cost:
$\mathcal{W}_{\text{updated}}=\underbrace{CA_{\text{former}}B_{\text{train}}R}_{\text{CABR-LoRA}}+\mathcal{W}_{\text{base}}$ (10)
M2: Frozen Base Merge
Preserves base weights using dual SECURA instances:
$\begin{split}\mathcal{W}_{\text{merged}}=&\underbrace{CA_{\text{frozen}}B_{\text{accum}}R}_{\text{Previous CABR-LoRA}}+\underbrace{CA_{\text{train}}B_{\text{reset}}R}_{\text{New CABR-LoRA}}\\ &\hskip 110.00017pt+\mathcal{W}_{\text{base}}\end{split}$ (11)
where parameter updates follow:
$A_{\text{former}},A_{\text{frozen}}$ $\leftarrow A_{\text{train}}$ (12)
$B_{\text{accum}}$ $\leftarrow B_{\text{accum}}+B_{\text{train}}$ (13)
[…]
M1 enables stronger fine-tuning by modifying base weights, while M2 prioritizes knowledge retention through:
  • Parameter isolation: Base weights remain frozen
  • Progressive accumulation: $B_{\text{accum}}$ aggregates historical updates
  • Cyclic resetting: $B_{\text{train}}$ reinitialization prevents update saturation
Our ablation study (4) reveals task-dependent superiority:
  • M1 for performance-critical domains (mathematics, programming)
  • M2 for knowledge-sensitive tasks (medical, economics)
The hybrid approach balances both objectives, with M1 generally providing better balanced optimization results.
Figure 3: Comparison of Weight SVD Norm across different methods: The "Basic" method represents the original LLM weights without fine-tuning. After 2000 training steps, DoRA shows a 0.517 decrease in norm, indicating some loss of information. SECURA preserves more knowledge with a minor 0.061 change, while LoRA shows a significant 3.076 increase, reflecting greater weight modification. Hyperparameter settings are detailed in Appendix D.
@@ 4.1 Experimental Setup
To evaluate catastrophic forgetting, we combined three QA datasets-ARC-Challenge(ARC-C) (Clark et al., 2018), GSM8K (Cobbe et al., 2021), Multipl-E (Cassano et al., 2022)-as the test set , which were used for training/fine-tuning the selected LLMs. This fusion dataset will be referred to as F-task in following parts. Further details can be found in the appendix C.
@@
Experiment 3 (Catastrophic Forgetting with Knowledge retention method): Benchmarks against continual learning methods at 1e-5 learning rate with sequential training on 16 tasks (each trained for 2K steps, totaling 32K steps) every PEFT based training method will fusion itself to frozen basic weights(e.g. CUR-LoRA, I-LoRA, SECURA) after one training task is done, measuring remain knowledge by using the F-task set and measured by BLEU, ROUGE-1, ROUGE-L.
Note: All main experiments’ SECURA uses the M1 merge method. For hyperparameter details, refer to the appendix D.